The processing of Q online orders can be done using hours
of robot time (denoted by K) and hours of human time
(denoted L). The processing of Q = 48 online orders requires either 4 hours of robot time (L = 0 and K = 4)
or 8 hours of human time (L = 8 and K = 0). Hours
of robot time and hours of human time are perfect substitutes; for example, the processing of Q = 48 online
orders could also be done using K = 3
Determine the labor and capital demand in terms
of Q0, w and r
1
Expert's answer
2020-03-20T09:47:53-0400

The labor or capital demand curves show the value of the marginal product of labor (MPL) or capital (MPK) respectively as a function of quantity of labor or capital hired.

MPL = 48/8 = 6 units,

MPK = 48/4 = 12 units.

Q = 6L + 12K.